# RadioActive: 3D Radiological Interactive Segmentation Benchmark

### v2

@@ Abstract
Current interactive segmentation approaches, inspired by the success of META’s Segment Anything model, have achieved notable advancements, however, they come with substantial limitations that hinder their practical application in 3D radiological scenarios. These include unrealistic human interaction requirements, such as slice-by-slice operations for 2D models on 3D data, a lack of iterative interactive refinement, and insufficient evaluation experiments. These shortcomings prevent accurate assessment of model performance and lead to inconsistent outcomes across studies. The RadioActive benchmark overcomes these challenges by offering a comprehensive and reproducible evaluation of interactive segmentation methods in realistic, clinically relevant scenarios. It includes diverse datasets, target structures, and interactive segmentation methods, and provides a flexible, extendable codebase that allows seamless integration of new models and prompting strategies. We also introduce advanced prompting techniques to enable 2D models on 3D data by reducing the needed number of interaction steps, enabling a fair comparison. We show that surprisingly the performance of slice-wise prompted approaches can match native 3D methods, despite the domain gap. Our findings challenge the current literature and highlight that models not specifically trained on medical data can outperform the current specialized medical methods. By open-sourcing RadioActive, we invite the research community to integrate their models and prompting techniques, ensuring continuous and transparent evaluation of interactive segmentation models in 3D medical imaging.
@@ 1 Introduction
Due to their great potential, interactive segmentation methods have attracted a lot of attention in the medical domain and led to many studies evaluating and adapting SAM for 3D medical image segmentation [34, 5, 15, 45, 30, 3, 24, 11]. Moreover, several researchers have been inspired by SAM’s capabilities to develop their own methods, often specifically designed for the 3D nature of radiological data [9, 12, 21, 39].
@@
Neglecting refinement (P2): Many studies assess interactive segmentation methods based on a single interaction step, overlooking the inherent ambiguities in radiological images [24, 9, 11, 2]. Often, a second interaction may be necessary to specify which specific substructure the clinician wants to segment. This could be, e.g. a vessel within the liver, or the necrosis within a tumor, as exemplified in the well-known BraTs segmentation challenge [4]. Furthermore, clinicians often want to adapt the segmentations to their clinic’s local protocol or refine them, particularly for targets with high inter-rater variability, like pathological structures [10, 1, 14]. Overall, there is a notable lack of research exploring realistic, iterative interaction methods for 2D models applied to 3D volumes.
Obfuscated and insufficient evaluation (P3): Currently, there is a lack of a standardized approach to evaluation. Often results are not incomparable between papers, with evaluation often obfuscated or insufficient. We observed the following shortcomings: (i) Not specifying whether predictions were interactively refined or based on a single prompt with multiple points [3, 39]. (ii) Being intransparent on the number of initial prompts given [9]. (iii) Using the best mask rather than the final mask after interactive refinement [39]. (iv) Evaluating predictions slice-by-slice or on sub-patches of a 3D volume instead of evaluating on the full image [34, 24, 3, 12, 21]. (v) Excluding targets considered too small, hence neglecting valid targets such as small lesions that are neither tested nor trained on [24, 3, 39]. (vi) Comparing against non-promptable models and SAM, rather than any other promptable model trained on medical data [3, 24, 11, 12]. (vii) Lastly, overemphasizing segmenting healthy structures, such as organs, where existing supervised public models already perform well [40, 37], instead of focusing on pathologies, where interactive refinement could provide the greatest benefits [39, 43].
To overcome these challenges, establishing a benchmark is essential, aligning with the recent recommendations by Marinov et al. [26]. To this end, we present RadioActive, a robust, reproducible, and extendable Radiological InterActive Benchmark for 3D medical image segmentation. This benchmark highlights the most effective 2D and 3D interactive segmentation approaches and tests realistic prompting methods in the 3D domain. We conduct evaluations directly in the 3D domain and across multiple models, providing more generalizing insights, as opposed to prior work [16, 26, 34]. Our key contributions include:
@@
Through open-sourcing RadioActive, we invite researchers to integrate their methods into our framework, promoting continuous and equitable assessment that allows us to track the overall progress in the field of interactive 3D medical image segmentation reproducibly and transparently.
Figure 3: RadioActive overview. Although our evaluation is performed on entire 3D volumes, the benchmark accommodates both 3D and 2D interactive segmentation methods. While 3D model prompting is relatively straightforward, we introduce prompting and refinement strategies for 2D models that minimize the effort required from human interaction. The benchmark is designed to be extensible, and researchers are encouraged to propose and integrate additional methods seamlessly using our codebase particularly for areas marked by three dots.
@@ 2 RadioActive
The RadioActive Benchmark is designed to easily enable a fair and reproducible evaluation of 2D and 3D interactive segmentation methods for 3D radiological image segmentation for the very first time Fig. 3. While prompting 3D models is generally straightforward, we introduce specific prompting and refinement strategies for 2D models to streamline human interaction and reduce the simulated Human Effort. The proposed benchmark includes seven established models and ten datasets covering different target structures and image modalities. All datasets are publicly available and we provide an automatic download and pre-processing for improved usability and reproducibility.
Moreover, the benchmark is built with flexibility in mind, enabling seamless integration of additional methods. Researchers are invited to contribute new approaches, particularly new models, new prompting schemes, and new interesting datasets to the collection. Overall, the design of our benchmark allows for easy testing and validation of novel segmentation methods, making the benchmark a catalyst for advancing methodology for interactive 3D medical image segmentation. In the following, we present the different components of RadioActive.
@@ 2.1 Initial Prompting
Figure 4: Different promoting schemes for 2D models based on point prompts (on the left) and box prompts (on the right). While a) and b) expect unrealistic human slice-by-slice interaction, c) and d) illustrate the proposed prompt interpolation schemes, where a human needs to provide prompts for at least 3 slices (4 slices in this case). Prompts for the remaining slices are generated by interpolating between the initial prompts. e) and f) present the proposed prompt propagation methods, where the prompt for each subsequent slice is automatically generated based on the model’s prediction from the previous slice. Only the initial slice and upper and lower boundaries require manual prompts.
@@ 2.2 Refinement Prompting
For 2D models, we then individually feed all slices $i\in I$ where the voxel along the scribble was not predicted, along with the new positive prompt derived from the scribble and the previous prediction $s\subset S$, back into the model. For 3D models, we feed the whole 3D patch, together with the previous prediction $S$, and all new positive points derived from the scribble into the network in one step.
If negative prompts are selected, we identify a non-axial slice $S_{fp}$ of $S$ that contains the most false positives. Then we generate a contour curve around the ground truth target object at a distance of 2 pixels. We then select a subpart $C$ with a length of 60% of the full curve and sample all pixels $c\in C$ that are false positives to obtain a set of points $D$, simulating a user drawing a few scribbles in areas where the model over-segmented the target. For 2D models, we then generate new slice predictions for each slice containing a point in $D$ by providing the model with the previous prediction as well as new negative prompts: all $d\in D$ which belong to that slice. For 3D models, we again feed the whole 3D patch, together with the previous prediction $S$ and a negative prompt sampled from $D$. The motivation for the different approaches for FPs is, that correcting the model near the target boundary is likely to provide more informative feedback. As none of the evaluated methods features a dedicated scribble prompt encoding, we instead generate a set of point prompts from our scribble refinement trajectory, which is similar to the recently proposed work of Wong et al. [41] for scribble prompting of SAM. Aside from this refinement strategy, we also support unrealistic refinement prompts for which details are provided in the appendix in Sec. 6.
@@ 2.4 Interactive Segmentation methods
SAM-Med 3D incorporates a transformer-based 3D image encoder, 3D prompt encoder, and 3D mask decoder. It was trained from scratch using 22,000 3D images and 143,000 corresponding 3D masks and supports point and mask prompts and also allows for refinement [39].
@@ 2.5 Datasets
Nevertheless, we assembled a diverse collection of ten lesser-known or recently released public datasets featuring various pathologies and organs, including CT and MRI image modalities. Specific details of these are provided in Tab. 1. While the quantity of cases may seem low, many contain multiple instances or target structures. Moreover, due to the pathological nature of many they represent the currently common size of datasets in the medical domain. To enhance reproducibility and eliminate barriers of entry for non-domain experts, we automated the dataset download and pre-processing, minimizing any required domain knowledge to use the benchmark. However, due to the sparsity of labeled datasets, we urge developers to exclude these datasets from their train dataset selection, as inclusion would compromise the integrity of a clean evaluation through RadioActive.
Table 1: Overview over all Datasets. None of these datasets were part of the original training data for the methods, except for SegVol, which utilized D2 HanSeg.
@@ 3.1 Initial prediction
As an upper baseline, we begin with an idealized and unrealistic scenario where each slice is prompted individually for all 2D models. In this setting, we evaluate different numbers of point prompts per slice (PPS), as well as alternating positive and negative prompts ($\pm$ PPS), and slice-wise box prompts with varying numbers of boxes per slice (BPS). Figure 5 shows that models employing box prompts achieved significantly higher average Dice scores, with SAM2 demonstrating the strongest performance across all models. Conversely, point-based prompts performed poorly, particularly for small target regions, such as small MS lesions in dataset D1 (see Table 2 in the appendix). Although including positive and negative prompts and increasing the number of point prompts led to improvements, these were minor compared to the marked superiority of box-based prompts. Interestingly, generalist models not trained with a particular focus on medical data performed worse with point prompts per slice but outperformed models trained on medical data when using box prompts. This may be because models trained with medical data are better at resolving the ambiguity of point prompts, especially given the often low contrast between pathologies and surrounding tissue. In contrast, SAM and SAM2 models excel in object identification when using box prompts.
Figure 5: Unrealistic prompting of 2D Boxes each slice performs best. 2D models prompted with one Box Prompt Per Slice (BPS), indicated with a star, outperform models prompted with various Point Prompts per Slice (PPS, line plots). Drawing a 2D box accounts for 2 interactions similar to providing 2 points. While alternating positive and negative points (dashed lines) is slightly superior to only positive points the gap between points and boxes remains large.
@@
    For all models, simple prompt interpolation methods achieve similar performance as the corresponding prompt per-slice methods. For example for SAM-Med 2D, 3 Point Interpolation even outperforms prompting 1 PPS having a mean DSC of 26.08 compared to a DSC of 23.13. This is because we select central points for the interpolation methods instead of random foreground points. Another example is SAM prompted using 5 Box Interpolation, that has a mean DSC of 56.9 compared to a DSC of 57.84 when prompted with a box for each slice.
@@
    Among 3D models, only SegVol is competitive to 2D models that use box prompts.
@@
    Both box and point propagation perform worse than their interpolation counterpart, though this may improve as models evolve.
Figure 6: Simple Interpolation Strategies Match Unrealistic Slice-Wise Prompting. Sampling prompts from the interpolated connection between three initial prompts yields similar performance for SAM2 as slice-wise prompting (left). This is also observed for box interpolation across all models (middle). 3D models perform worse than 2D methods when only a few points are provided, while SegVol demonstrates that using a 3D box is superior to points (right).
3.2 Interactive Refinement
Finally, we evaluate the performance of the models during iterative refinement. For 2D models, this involves predicting on a slice-by-slice basis. As illustrated in Fig. 7 (top), adding refinement prompts to each slice results in a substantial performance boost. While the proposed scribble-based refinement consistently improves outcomes, it does not achieve the same level of improvement as adding a prompt to every slice, which is expected since not all imperfect slices are guaranteed to receive new prompts during the scribble refinements. Moreover, we observed that for 2D models, it is crucial to provide the initial prompts again for each of the refinement steps. In the appendix in Table 5, we present results without including the previous point in the iterative refinement step, which resulted in a performance decline during refinement. 2D models tend to over-segment the target, filling the entire slice foreground. The absence of the initial prompt leads to a complete loss of target location information, as the initial predicted mask is highly inaccurate. Our refinement likely generates negative additional prompts due to the large number of false positive pixels. Qualitative examples of this failure case are presented in the appendix in Sec. 7 in the appendix.
For 3D models, iterative refinement also led to consistent performance improvements. Both randomly sampled prompts and those derived from refinement scribbles improved performance with each refinement iteration. Although SegVol initially performs best for the initial prediction and achieves a mean DSC of 43.32 prompted with 5 centered point prompts, it lacks support for further refinement. In contrast, SAM-Med 3D Turbo - worse in the initial prediction with a Dice of 40.89 - surpasses SegVol after several refinement steps reaching a DSC of 47.4.
Figure 7: All models demonstrate significant improvements from iterative refinement. Dashed lines represent the use of the proposed scribble refinement. While the unrealistic scenario of one refinement point per slice yields better performance, the proposed scribble refinement consistently enhances results across iterations for 2D models.

### v3

@@ Abstract
Effortless and precise segmentation with minimal clinician effort could greatly streamline clinical workflows. Recent interactive segmentation models, inspired by META’s Segment Anything, have made significant progress but face critical limitations in 3D radiology. These include impractical human interaction requirements such as slice-by-slice operations for 2D models on 3D data and a lack of iterative refinement. Prior studies have been hindered by inadequate evaluation protocols, resulting in unreliable performance assessments and inconsistent findings across studies. The RadioActive benchmark addresses these challenges by providing a rigorous and reproducible evaluation framework for interactive segmentation methods in clinically relevant scenarios. It features diverse datasets, a wide range of target structures, and the most impactful 2D and 3D interactive segmentation methods, all within a flexible and extensible codebase. We also introduce advanced prompting techniques that reduce interaction steps, enabling fair comparisons between 2D and 3D models. Surprisingly, SAM 2 outperforms all specialized medical 2D and 3D models in a setting requiring only a few interactions to generate prompts for a 3D volume. This challenges prevailing assumptions and demonstrates that general-purpose models surpass specialized medical approaches. By open-sourcing RadioActive, we invite researchers to integrate their models and prompting techniques, ensuring continuous and transparent evaluation of 3D medical interactive models.
@@ 1 Introduction
Figure 2: RadioActive overview. Although our evaluation is performed on entire 3D volumes, the benchmark accommodates both 3D and 2D interactive segmentation methods. While 3D model prompting is relatively straightforward, we introduce prompting and refinement strategies for 2D models that minimize the effort required from human interaction. The benchmark is designed to be extensible, and researchers are encouraged to propose and integrate additional methods seamlessly using our codebase, particularly for areas marked by three dots.
@@
Due to their great potential, interactive segmentation methods have attracted a lot of attention in the medical domain and led to many studies evaluating and adapting SAM for 3D medical image segmentation [38, 7, 18, 50, 33, 4, 27, 13, 45]. Moreover, several researchers have been inspired by SAM’s capabilities to develop their own methods, often specifically designed for 3D medical data [11, 15, 24, 43].
@@
Neglecting refinement (P2): Many studies assess interactive segmentation methods based on a single interaction step, overlooking the inherent ambiguities in radiological images [27, 11, 13, 3]. Often, a second interaction may be necessary to specify which specific substructure the clinician wants to segment. This could be, e.g. a vessel within the liver, or the necrosis within a tumor, as exemplified in the well-known BraTs segmentation challenge [6]. Furthermore, clinicians often want to adapt the segmentations to their clinic’s local protocol or refine them, particularly for targets with high inter-rater variability, like pathological structures [12, 2, 17]. Overall, there is a notable lack of research exploring realistic, iterative refinement methods for 2D models applied to 3D volumes.
Obfuscated and insufficient evaluation (P3):Currently, no standardized evaluation framework exists, leading to inconsistencies across studies. As a result, comparisons between papers are often unreliable, with evaluations frequently being opaque or insufficient. We observed the following shortcomings: (i) Not specifying whether predictions were interactively refined or based on a single prompt with multiple points [4, 43]. (ii) Being intransparent on the number of initial prompts given [11]. (iii) Using the best mask rather than the final mask after interactive refinement [43]. (iv) Evaluating predictions slice-by-slice or on sub-patches of a 3D volume instead of evaluating the full image [38, 27, 4, 15, 24]. (v) Excluding targets considered too small, hence neglecting valid targets such as small lesions that are neither tested nor trained on [27, 4, 43]. (vi) Comparing against non-promptable models and SAM, rather than any other promptable model trained on medical data [4, 27, 13, 15]. (vii) Lastly, overemphasizing segmenting healthy structures, such as organs, where existing supervised public models already perform well [44, 41], instead of focusing on pathologies, where interactive refinement could provide the greatest benefits [43, 47].
Developing a benchmark is key to overcoming these challenges, in line with the recent recommendations by Marinov et al. [29]. To this end, we present RadioActive, a robust, reproducible, and extendable Radiological InterActive Benchmark for 3D medical image segmentation. This benchmark highlights the most effective 2D and 3D interactive segmentation approaches and tests realistic prompting methods in the 3D domain. We conduct evaluations directly in the 3D domain and across multiple models and datasets. Our key contributions include:
@@
Through open-sourcing RadioActive, we invite researchers to integrate their methods into our framework, promoting continuous and equitable assessment that allows reproducibility and transparently tracking the overall progress in interactive 3D medical image segmentation.
@@ 2 RadioActive
The RadioActive Benchmark is designed to easily enable a fair and reproducible evaluation of 2D and 3D interactive segmentation methods for 3D radiological image segmentation for the very first time Fig. 2. In Sec. 6, we include an in-depth overview of the state of current independent benchmarkings for interactive segmentation methods for 3D radiological images [19, 29, 38, 30, 14, 5, 48, 1]. While prompting 3D models is generally straightforward, we introduce specific prompting and refinement strategies for 2D models to streamline human interaction and reduce the simulated Human Effort. The proposed benchmark includes seven established models and ten datasets covering different target structures and image modalities. All datasets are publicly available and we provide an automatic download and pre-processing for improved usability and reproducibility. Moreover, the benchmark is built with flexibility in mind, enabling seamless integration of additional methods. Researchers are invited to contribute new approaches, particularly new models, new prompting schemes, and new interesting datasets to the collection. Overall, the design of our benchmark allows for easy testing and validation of novel segmentation methods, making the benchmark a catalyst for advancing methodology for interactive 3D medical image segmentation. In the following, we present the different components of RadioActive.
@@ 2.1 Initial Prompting
Figure 4: Different prompting schemes for 2D models based on point prompts (on the left) and box prompts (on the right). While a) and b) expect unrealistic human slice-by-slice interaction, c) and d) illustrate the proposed prompt interpolation schemes, where a human needs to provide prompts for at least 3 slices (4 slices in this case). Prompts for the remaining slices are generated by interpolating between the initial prompts. e) and f) present the proposed prompt propagation methods, where the prompt for each subsequent slice is automatically generated based on the model’s prediction from the previous slice. Only the initial slice and upper and lower boundaries require manual prompts.
@@ 2.2 Refinement Prompting
For 2D models, we then individually feed all slices $i\in I$ where the voxel along the scribble was not predicted, along with the new positive prompt derived from the scribble and the previous prediction $s\subset S$, back into the model. For 3D models, we feed the whole 3D patch, together with the previous prediction $S$, and a random point derived from the scribble into the network in one step.
If negative prompts are selected, we identify a non-axial slice $S_{fp}$ of $S$ that contains the most false positives. Then we generate a contour curve around the ground truth target object at a distance of 2 pixels. We then select a subpart $C$ with a length of 60% of the full curve and sample all pixels $c\in C$ that are false positives to obtain a set of points $D$, simulating a user drawing a few scribbles in areas where the model over-segmented the target. For 2D models, we then generate new slice predictions for each slice containing a point in $D$ by providing the model with the previous prediction as well as new negative prompts: all $d\in D$ which belong to that slice. For 3D models, we again feed the whole 3D patch, together with the previous prediction $S$ and a negative prompt sampled from $D$. The motivation for the different approaches for FPs is, that correcting the model near the target boundary is likely to provide more informative feedback. A visualization of the proposed scribble refinement is provided in the appendix in Fig. 9. As none of the evaluated methods features a dedicated scribble prompt encoding, we instead generate a set of point prompts from our scribble refinement trajectory, which is similar to the recently proposed work of Wong et al. [45] for scribble prompting of SAM. Aside from this refinement strategy, we also support unrealistic refinement prompts for which details are provided in the appendix in Sec. 7.
@@ 2.4 Interactive Segmentation methods
SAM-Med 3D incorporates a transformer-based 3D image encoder, 3D prompt encoder, and 3D mask decoder. It was trained from scratch using 22,000 3D images and 143,000 corresponding 3D masks and supports point and mask prompts and also allows for refinement [43]. Notably, it supports only one point prompt per prediction.
@@
ScribblePrompt is an interactive 2D segmentation model designed specifically for medical imaging [45]. It supports multiple input prompts, including points, boxes, and masks, while introducing scribbles as a novel interaction method. The model integrates these prompts as an additional channel within a U-Net architecture and is trained with 54,000 scans.
@@ 2.5 Datasets
Nevertheless, we assembled a diverse collection of ten lesser-known or recently released public datasets featuring various pathologies and organs, including CT and MRI image modalities. Specific details of these are provided in Tab. 1. While the quantity of cases may seem low, many contain multiple instances or target structures. Moreover, due to the pathological nature of many they represent the currently common size of datasets in the medical domain. Overall, these 10 datasets offer a robust foundation for evaluating interactive segmentation methods, ensuring diverse, clinically relevant challenges while mitigating biases from training data overlap. To enhance reproducibility and eliminate barriers of entry for non-domain experts, we automated the dataset download and pre-processing, minimizing any required domain knowledge to use the benchmark. However, due to the sparsity of labeled datasets, we urge developers to exclude these datasets from their train dataset selection, as inclusion would compromise the integrity of a clean evaluation through RadioActive.
Table 1: Overview over all Datasets. None of these datasets were part of the original training data for the methods, except for SegVol, which utilized CT images from D2 HanSeg .
@@ 3.1 Initial prediction
As an upper baseline, we begin with an idealized and unrealistic scenario where each slice is prompted individually for all 2D models. In this setting, we evaluate different numbers of point prompts per slice (PPS), as well as alternating positive and negative prompts ($\pm$ PPS), and slice-wise box prompts with one box per slice (BPS). Figure 5 shows that models employing box prompts achieved significantly higher average Dice scores, with SAM2 demonstrating the strongest performance across all models. Conversely, point-based prompts performed poorly, particularly for small target regions, such as small MS lesions in dataset D1 (see Tab. 2 in the appendix). Although including positive and negative prompts and increasing the number of point prompts led to improvements, only ScribblePrompter achieved performance in the same range as box-based prompts. Interestingly, generalist models not trained with a particular focus on medical data performed worse with point prompts per slice but outperformed models trained on medical data when using box prompts. This may be because models trained with medical data are better at resolving the ambiguity of point prompts, especially given the often low contrast between pathologies and surrounding tissue.
Figure 5: Unrealistic prompting of 2D Boxes each slice performs best. 2D models prompted with one Box Prompt Per Slice (BPS), indicated with a star, outperform models prompted with various Point Prompts per Slice (PPS, line plots). Drawing a 2D box accounts for 2 interactions similar to providing 2 points. Providing alternating positive and negative points (dashed lines) is slightly superior to only positive points.
@@
    For all models, simple prompt interpolation methods achieve similar performance as the corresponding prompt per-slice methods. SAM2 achieves a mean Dice of 81.88 when we using 5 initial 2D boxes (5 box inter) compared to a Dice of 83.63 for one box per slice (Box PS), as can be seen in the appendix in Tab. 3.
@@
    Among 3D models, only SegVol is competitive to 2D models that use box prompts or 5-10 points per volume as can be seen in the appendix in Tab. 4.
@@
    Both box and point propagation perform worse than their interpolation counterpart. For propagation schemes, if a previous prompt did not yield a prediction, the propagation would terminate. However, in practice, we observed the opposite effect. Notably, point prompts frequently caused severe oversegmentation, covering the entire body and obscuring the target object, ultimately leading to a loss of meaningful information. Qualitative examples of this failure case are presented in the appendix in Sec. 5. This issue may improve as models continue to evolve.
Figure 6: Simple Interpolation Strategies Match Unrealistic Slice-Wise Prompting. Sampling prompts from the interpolated connection between three initial prompts yields similar performance for SAM2 as slice-wise prompting (left). This is also observed for box interpolation across all models (middle). 3D models perform worse than 2D methods. We compare points sampled at equal axial distances along the five-point interpolation prompter line, which approximates the axial target centerline, with randomly sampled points. Sampling a single point from the center performs better than sampling a random point, but random foreground sampling is better for multiple points.
3.2 Iteractive Refinement
Finally, we evaluate the performance of the models during iterative refinement. For 2D models, this involves predicting on a slice-by-slice basis. As illustrated in Fig. 7, adding refinement prompts to each slice results in a substantial performance boost. While the proposed scribble-based refinement consistently improves outcomes, it does not achieve the same level of improvement as adding a prompt to every slice, which is expected since not all imperfect slices are guaranteed to receive new prompts during the scribble refinements. Notably, SAM2 improves more compared to SAM. We observed that for 2D models, it is crucial to provide the initial prompts again for each of the refinement steps. In the appendix in Tab. 5, we present results without including the previous point in the iterative refinement step, which resulted in a performance decline during refinement. As demonstrated in our previous experiments, point prompts do not produce reliable initial segmentations. Although their refinement consistently enhances results, overall performance remains suboptimal. To overcome this, we introduce a new baseline for future interactive segmentation methods: initializing 2D models with three interpolated 2D boxes, followed by Scribble refinement. This approach achieves the best overall performance with Sam2.
For 3D models, iterative refinement also led to consistent performance improvements. Both randomly sampled prompts and those derived from refinement scribbles improved performance similarly. However, they do not reach similar performance as 2D models.
Figure 7: Iterative refinement is essential. All methods improve using the proposed scribble refinement (dashed) or using a random refinement point per volume or slice (solid line). When given a strong initial prediction, as with our proposed 3-box interpolation prompting, 2D models achieve superior performance.